OpenStudy (anonymous):

Solve using substitution method x+y=6 y=x-4

OpenStudy (anonymous):

x + (x-4)) = 6 2x - 4 = 6 2x = 10 x = 5 Plug that back in y = 5-4 y = 1 Answer is (5,1)